Suppose that someone offers you a bet concerning the tossing of a pair of fair dice. If both die turn up the same number you will win $240. Otherwise you must pay $60. Your expected winning is ?
1
Expert's answer
2020-08-02T15:34:22-0400

The probability of both die turn up the same number is 6/36 = 1/6.

Your expected winning is:

"240\u00d71\/6 + (-60)\u00d7(1 - 1\/6) = - 10."